OpenStudy (anonymous):

The graph of f passes through (-1,3) and is perpendicular to the line that has an x-intercept of 5 and a y-intercept of -40. The slope-intercept form of y=? I attempted to answer this with 6/43x+135/43, but it was marked wrong.

OpenStudy (theprofessor35):

So, lets think.... Slope intercept form is y = mx + b

OpenStudy (theprofessor35):

Find the slope using the equation \[m = \frac{ y2 - y1 }{ x2 - x1 }\]

OpenStudy (anonymous):

It would be -43/6.

OpenStudy (anonymous):

And switching it to 6/43 would make the slope perpendicular.

OpenStudy (theprofessor35):

Hmmm, good question....this is quite confusing

OpenStudy (anonymous):

Yep, it is. I don't know where I made a mistake.

OpenStudy (priyar):

first write the eq in intercept form

OpenStudy (priyar):

like: x/a +y/b = 1 where a=5 and b=-40. what do u get?

OpenStudy (priyar):

@staysmart ?

OpenStudy (triciaal):